Given the function f(x) = 2x + 5. If the domain is {2,3,4,5} list the ordered pairs to describe the function. Please show how you arrived at your answer.
Answer by Flannery(124)

f(x) is what x becomes when plugged into 2x+5.
they gave you four values for x.
domain values are x values and range values are y values (y may be described as f(x) if and only if y is a function of x).
y is a function of x if and only if for every value x the operations yeild one and only one output y.
2(2)+5=9 (2,9)
2(3)+5=11 (3,11)
2(4)+5=13 (4,13)
2(5)+5=15 (5,15)
ordered pairs are listed (x,y)
f(x)=2x+5 is a linear equation in slope intercept form.
the slope is 2 and the y intercept is at (0,5).